The average density of a white dwarf of radius 109 cm is 106 gm/cm3. Is general theory of relativity needed to study the dynamics of this star? What happens if the star shrinks to a radius thousand times smaller?
1
Expert's answer
2019-04-01T10:33:46-0400

The general theory of relativity for such an object becomes important when its physical radius "R" becomes comparable to its Schwarzschild radius, which is given by the formula

"r_s = \frac{2 G M}{c^2} \, ,"

where "G" is the Newton's gravitational constant, "M" is the mass of the star, and "c" is the speed of light. Estimating the mass as "M = 4 \pi \rho R^3 / 3", where "\rho" is the average density, we conclude that the importance of the general theory of relativity is determined by the magnitude of the ratio

"\frac{r_s}{R} = \frac{8 \pi G \rho R^2}{3 c^2} \, ."

It remains to substitute the values "G = 6.67 \times 10^{-11}\,\text{m}^3 / \text{kg s}^2", "\rho = 10^6\, \text{g} / \text{cm}^3 = 10^9\, \text{kg} / \text{m}^3", "R = 10^9\, \text{cm} = 10^7\, \text{m}", and "c = 3 \times 10^8\, \text{m} / \text{s}". We obtain

"\frac{r_s}{R} \approx 6 \times 10^{-4} \, ."

The smallness of this ratio means that the effects of general relativity are insignificant. If the star shrinks to a radius thousand times smaller, its mass and, therefore, its Shwarzschild radius remain the same. Hence, the ratio "r_s\/R" increases thousand times and reaches the value of about 0.6. In this case, the general theory of relativity will be needed to study the dynamics of the star.